OpenStudy (anonymous):

Find a polynomial f (x) that satisfies the equation xf''(x)+f(x)=x^2

OpenStudy (anonymous):

I'd start by looking for the minimal degree polynomial. You can check for yourself that a polynomial of degree less than 2 won't work. \[f(x)=a_0+a_1x+a_2x^2\\ f'(x)=a_1+2a_2x\\ f''(x)=2a_2\] Subbing this into the equation gives \[2a_2x+a_0+a_1x+a_2x^2=x^2\] You have to satisfy a few conditions here: \[\begin{cases} a_2=1&\text{match up coefficients of quadratic terms}\\ a_1+2a_2=0&\text{match up linear coefficients}\\ a_0=0&\text{constant coefficients} \end{cases}\] This gives \(a_0=0\), \(a_1=-2\), and \(a_2=1\), so one possible polynomial solution is \[f(x)=x^2-2x\]